THE FIXED POINT PROPERTY FOR POSETS OF SMALL WIDTH

JONATHAN DAVID FARLEY

Abstract. The fixed point property for finite posets of width 3 and 4 is studied in terms of forbidden retracts. The ranked forbidden retracts for width 3 and 4 are determined explicitly. The ranked forbidden retracts for the width 3 case that are linearly indecomposable are examined to see which are minimal automorphic. Part of a problem of Niederle from 1989 is thus solved.